Determine if the system has a nontrivial solution. Try to use as few row operations as possible 4x1-6x2 + 13x3 = 0 -4x1-2x2-7x3 = 0 8x1 4x2+14x30 Choose the correct answer below. O A. O B. O C. The system has a nontrivial solution. It is impossible to determine. The system has only a trivial solution.
Answer #1

system is

4 -2 -7r20

augmented matrix is

4 -6 13
-4 -2 -7
8 4 14

convert into Reduced Row Eschelon Form...

Divide row1 by 4

1 -3/2 13/4
-4 -2 -7
8 4 14


Add (4 * row1) to row2

1 -3/2 13/4
0 -8 6
8 4 14


Add (-8 * row1) to row3

1 -3/2 13/4
0 -8 6
0 16 -12


Divide row2 by -8

1 -3/2 13/4
0 1 -3/4
0 16 -12


Add (-16 * row2) to row3

1 -3/2 13/4
0 1 -3/4
0 0 0


Add (3/2 * row2) to row1

1 0 17/8
0 1 -3/4
0 0 0

reduced system is

\begin{pmatrix}1&0&\frac{17}{8}\\ 0&1&-\frac{3}{4}\\ 0&0&0\end{pmatrix}\begin{pmatrix}x_1\\ x_2\\ x_3\end{pmatrix}=0

x_1+\frac{17}{8}x_3=0.........{\color{Red} x_1=-\frac{17}{8}x_3}

x_2-\frac{3}{4}x_3=0.........{\color{Red} x_2=\frac{3}{4}x_3}

{\color{Red} x_3=x_3}.................free variable

non trivial solution is

{\color{Red} \begin{pmatrix}x_1\\ x_2\\ x_3\end{pmatrix}=\begin{pmatrix}-\frac{17}{8}\\ \frac{3}{4}\\ 1\end{pmatrix}x_3}